OpenStudy (karim728):

Fun Integral:

OpenStudy (karim728):

\[\int\limits_{}^{}\sqrt{1+\sin 2x } dx\]

ganeshie8 (ganeshie8):

\[\sqrt{\color{red}{1}+\sin (2x)} \\~\\=\sqrt{\color{red}{\sin^2x+\cos^2x}+\sin (2x)} \\~\\=\sqrt{(\sin x + \cos x)^2} \\~\\= |\sin x +\cos x|\]

OpenStudy (karim728):

dang have you seen this before?

ganeshie8 (ganeshie8):

I guess several time haha but what can be done with that absolute value ?

OpenStudy (karim728):

that absolute value is just like when you want to integrate the absolute value of x, you separate into the integral of x for all numbers equal or greater than 0 and the integral of -x for all numbers equal or less than 0